OpenStudy (anonymous):

Find the midpoint of a line segment joining the points P1 and P2. P1= (1,4) p2= (-5,6) is the midpoint answer (2,5) ??

OpenStudy (anonymous):

\[\frac{x_1+x_2}{2},\frac{y_1+y_2}{2}\]

OpenStudy (anonymous):

M= (x1+x2/2 , y1+y2/2) ?

OpenStudy (anonymous):

-2,5 it is

OpenStudy (anonymous):

oh yeah cause it was -4 ! thanks.